Not sure if it comes under this topic but I have tried to build up an equation to solve it and I think I can but I'm not understanding how. So here's the question :
In a picnic there were 240 persons. There were 20 more men than women and 20 more adults than children. How many men were there in the picnic ? Answer given in my book shows it's 75, but I dont understand how.

Answer by ikleyn(52781)

.
Not sure if it comes under this topic but I have tried to build up an equation to solve it and I think I can
but I'm not understanding how. So here's the question :
In a picnic there were 240 persons. There were 20 more men than women and 20 more adults than children.
How many men were there in the picnic ? Answer given in my book shows it's 75, but I dont understand how.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

M for men, W for women and C for children.

M + W + C = 240,  (1)
M - W     =  20,  (2)
M + W - C =  20.  (3).

Add equations (1) and (3). You will get

2*(M + W) = 260,  or

M + W = 130.      (4)

Add (2) and (4). You will get

2M = 150.

Hence, M = 75.

Solved.